If a body of mass 3 kg is dropped downwards, after 1 sec another ball is dropped downwards from the same point. What is the distance between them after 3 sec.

(a) 25 m                (b) 20 m

(c) 50 m                (d) 9.8 m

 

how do you solve this question?

Grade:12

1 Answers

Archita
8 years ago
for the 1st ball
distance moved in 3 seconds,
 
s=ut+(1/2)gt2
for a freely falling body u=0
 
s=(1/2)gt2
s=(1/2)×10×32
s=45m
 
for the 2nd ball
when real time = 3s the body has been let into free fall for only 2s as the 2nd ball was dropped one second after the first ball
hence t=2s
 
s=ut + (1/2)gt2
s=(1/2)×10×22
s=20m
 
distance between them =
 
                                             (45-20)m
                                            =25m
 
if u have any doubts ask in answer box.
approve if helpful